Wilbur left the hardware store and traveled toward the recycling plant at an average speed of 33km/h. Mary left 2 hours later and traveled in the same direction but with an avg. spd. of 55km/h. How long did Wilbuer travel before Mary caught up?
When I solved this, I got the time that Wilbur traveled before Mary caught up to him. But I am not sure if the phrase "How long did Wilbuer travel before Mary caught up?" is asking for the distance or the time. Any answers?

Answer by ankor@dixie-net.com(22740)

Wilbur left the hardware store and traveled toward the recycling plant at an average speed of 33km/h.
Mary left 2 hours later and traveled in the same direction but with an avg. speed of 55km/h.
How long did Wilbur travel before Mary caught up?
:
Let t = W's travel time
then
(t-2) = M's travel time
:
When M catches W they will have traveled the same distance from the store.
Write a distance equation
:
M's dist = W's dist
55(t-2) = 33t
55t - 110 = 33t
55t - 33t = 110
22t = 110
t = 110/22
t = 5 hrs W's travel time when M caught up; which is what they are asking for
:
:
We can check this by finding the distance each traveled, they should be the same
55(5-2) = 165 km
33(5) = 165 km